OpenStudy (jaylokss):

I'LL GIVE MEDAL What is the arc length of the arc subtended in a circle with radius 6 and an angle of 7pi/8

OpenStudy (luffingsails):

How many radians make up a complete circle?

OpenStudy (moldybubblegum12):

(21pi)/4 Because, from the formula s = rθ where s represents the arc length, θ represents the central angle in radians and r is the length of the radius. s = 6 x (7pi)/8 = 42π / 8 = 21π / 4